### v1

# Quantum reflection time and the Goos-Hänchen effect

Abstract
We explore the analogy between following the motion of a reflected wave packet, and a quantum measurement of the spatial delay imposed on the particle by the scattering potential. It is shown that converting such delays into temporal durations can lead to “negative times”, and best be avoided. It is also demonstrated that a “soft” potential can be replaced by superposition of hard walls. This representation is used for calculating the Goos-Hänchen shift of a Gaussian beam incident on a potential step or a barrier at an oblique angle.
I Introduction
The Goos-Hänchen (GH) effect, first observed in 1947 for electromagnetic waves GH0 , consists in lateral displacement of a beam reflected off a plane surface. The effect is not unique to electromagnetism, and can occur in a variety of wave phenomena, including quantum scattering, see for example GH2 -GH4 . The displacement is often related with the time delay, associated with the scattering process GH1 , and can be seen as a way of measuring it. However, there is still little consensus about the definition of the quantum mechanical time delay. Much attention has been paid to the “tunnelling time problem” REV1 -REV2 of defining the duration a tunnelling particle is supposed to spend in the barrier region. Recently we argued that the answer is provided by the measurement theory, in a uniquely quantum way DS1 , DS2 , DS3 . According to the Uncertainty Principle FeynL , to know the duration of interest one needs to destroy the interference, and that destroys tunnelling. Conversely, one can preserve tunnelling, but with interference intact the duration of interest must remain indeterminate DS1 , DS2 . The same should be true in quantum reflection, where one would expect similar non-classical behaviour for a particle repulsed by a potential well, or reflected above a potential barrier.
In this work, we revisit the quantum Goos-Hänchen effect in order to establish whether determination of the GH shift does amount to measuring the time spent by the particle in the scattering potential. We will answer the question by the negative, and proceed to determine what is being measured, and to what accuracy. The paper is organised as follows. In Section II, we will rewrite the reflected wave packet (WP) in terms of the freely propagating counterpart, and a number of its mirror images. In section III we exploit the analogy with a von Neumann quantum measurement performed on a pre- and post-selected quantum system, and identify the measured quantity. In Sect. IV we show that in the classical limit one does determine the duration spent by the classical trajectory in the scattering potential. In Sect. V we attempt to extend the idea to the fully quantum case, and fail. In Sect. VI obtain the reflection “phase time” and discuss its properties. Sects. VII and VIII offer two simple examples of what has been said so far. In Sect. IX we abandon the attempts to define a meaningful “reflection time”, and replace instead a soft potential barrier by a superposition of virtual hard walls. In Sect. X, we use this representation to reconstruct the reflected beam, and evaluate the GH shift in terms of the spatial shifts experienced by the reflected particle. Section XI contains our conclusions.
II Reflection in the momentum and coordinate representations
We consider a potential potential $V(x)$, contained in the region $0<x<d$, and a wave packet (WP) state $|\Psi(0){\rangle}$, ${\langle}x|\Psi(0){\rangle}\equiv 0$ for $x>0$.
Figure 1: a) A wave packet with a mean momentum $p$, incident on a potential barrier $V(x)$, is partly reflected. Its mean energy is lower than the barrier height, and $x_{<}$ denotes the classical turning point $V(x_{<})=E(p)$. b) Freely propagating WP and its reflection is a mirror placed at $x=x^{\prime}/2$. c) The reflected WP shown in (a) is a superposition of the WPs reflected by a variety of infinite walls.
A convenient choice is a Gaussian WP moving towards the barrier (see below). The initial state at $t=0$ can be expanded in plane waves, (we use $\hbar=1$)
$\Psi(x,0)=\int_{-\infty}^{\infty}dk{\langle}x|k{\rangle}{\langle}k|\Psi(0){\rangle},$ (1)
where ${\langle}x|k{\rangle}=(2\pi)^{-1/2}\exp(ikx)$. Time evolution of the WP is found by expanding it in the complete set of scattering states, incident on the barrier from the left and from the right (see Appendix A). Thus, for $t>0$ the wave function in the left half-space $x<0$ is given by
$\Psi(x,t)=\Psi_{0}(x,t)+\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$ (2)
$\int_{-\infty}^{\infty}\frac{dk}{\sqrt{2\pi}}R(k){\langle}k|\Psi(0){\rangle}\exp[-ikx-iE(k)t],\ \ [x<0],$
where
$\Psi_{0}(x,t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}\exp[ikx-iE(k)t],$ (3)
is the freely evolving initial state, and $R(k)$ is the reflection amplitude for a particle with momentum $k$, incident on the potential from the left. The function $R(k)$ is known to be analytic in the complex $k$-plane, with poles in its lower half, and also on the positive imaginary axis, if $V(x)$ supports bound states BZP . Note that Eq.(2) is valid even when if the initial state (1) contains negative momenta, in which case one uses $R(-|k|)=R^{*}(|k|)$ (see Appendix A).
One can transform Eq.(2) to the coordinate representation by inserting the identity
$R(k)=\int_{-\infty}^{\infty}dx^{\prime}\exp(ikx^{\prime})\xi(x^{\prime})$ (4)
$\xi(x^{\prime})=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\exp(-ik^{\prime}x^{\prime})R(k^{\prime})$ (5)
into Eq.(3). This yields
$\Psi(x,t)=\Psi_{0}(x,t)+\int_{-\infty}^{\infty}dx^{\prime}\Psi_{0}(x^{\prime}-x,t)\xi(x^{\prime}).$ (6)
where $\Psi_{0}(x^{\prime}-x,t)$ is the image of the freely propagated WP (3) in a mirror placed at $x=x^{\prime}/2$ (see Fig.1b). Each image enters the superposition (6) with a weight $\xi(x^{\prime})$, the Fourier transform of the reflection amplitude with respect to particle’s momentum. For a potential $V(x)$, such that $R(k)\to 0$ as $|k|\to\infty$, $\xi(x^{\prime})$ is a smooth function of $x^{\prime}$, and, in the absence of bound states, one has
$\xi(x^{\prime})\equiv 0\quad\text{for}\quad x^{\prime}<0.$ (7)
The evolution of the wave function in the left half-space can now be visualised as follows. At first, the WP localised far to the left, moves freely towards the scatterer, and the wave function is given by the first term in Eq.(2), $\Psi(x,t)\approx\Psi_{0}(x,t)$. For a while its mirror images $\Psi_{0}(x^{\prime}-x,t)$, moving in the opposite direction, remain “off screen” in the right half-space $x>0$. As the free WP crosses the origin, and disappears from the screen, its mirror images enter the region $x<0$, replacing there $\Psi_{0}(x,t)$. The scattering is finished when the left half-space contains only the reflected (R) part of the wave function
$\Psi(x,t){\xrightarrow[t\to\infty]{}}\int_{-\infty}^{\infty}dx^{\prime}\xi(x^{\prime})\Psi_{0}(x^{\prime}-x,t)\equiv\Psi_{R}(x,t).\quad\quad$ (8)
Needless to say, the representation (6) has no physical significance in the right region $x>0$, but is exact for any $x<0$.
III Analogy with a quantum measurement
Let the initial state of a particle of mass $m$ be a Gaussian WP with a mean momentum $p=mv>0$, and coordinate width $\Delta x$, placed a $x=x_{0}<0$. Throughout the paper we will neglect the WP’s spreading, so the free state in Eq.(3) propagates without distortion, (see Appendix B), and we have
$\Psi_{0}(x,t)\approx G_{0}(x-x_{0}-vt)\exp\left[ip(x-x_{0})-iE(p)t\right],$ (9)
where
$G_{0}(x)=\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left(-\frac{x^{2}}{\Delta x^{2}}\right).$ (10)
Reflecting $\Psi_{0}(x,t)$ about the point $x^{\prime}/2$ requires changing $x\to x^{\prime}-x$, and for the reflected state (8) we find
$\Psi_{R}(x,t)\approx\exp[-ip(x+x_{0})-iE(p)t]\times\quad\quad\quad\quad\quad$ (11)
$\int_{-\infty}^{\infty}dx^{\prime}G_{0}(x^{\prime}-X(t))\exp(ipx^{\prime})\xi(x^{\prime}),\quad[x<0],$
where we introduced a shorthand
$X(t)\equiv x+x_{0}+vt.$ (12)
Similarity with a quantum measurement made on a pre- and post-selected quantum system is now evident. The amplitude of finding the particle at $x$, $\Psi_{R}(x,t)$, matches that of a von Neumann pointer vN of a mass $m$ and momentum p-p, set up to determine the manner in which a post-selected system reaches its final state [see Eq.(79) of the Appendix D]. As with the von Neumann pointer, the amplitude $\Psi_{R}(x,t)$ can be seen as a result of interference between various alternatives, each labelled by $x^{\prime}$. The experimenter, who finds the reflected particle in $x$ at $t$, knows that $x^{\prime}$ (whatever it may be) lies in the interval $X(t)-\Delta x\lesssim x^{\prime}\lesssim X(t)+\Delta x$, since only the values within this range contribute to the corresponding probability $|\Psi_{R}(x,t)|^{2}$. Destruction of interference between the alternatives and, therefore, the accuracy with which $x^{\prime}$ is determined, is the greater the smaller is the WP’s width $\Delta x$.
The peculiarity of the situation is that, unlike in the von Neumann example, no other degree of freedom is involved, so the particle appears to “measure itself”. But what precisely is being measured? Rewriting the reflected plane wave as
$R(p)\frac{\exp(-ipx)}{\sqrt{2\pi}}=\int_{-\infty}^{\infty}\frac{dx^{\prime}}{\sqrt{2\pi}}\exp[-ip(x-x^{\prime})]\xi(x^{\prime}),\quad$ (13)
we note that $x^{\prime}$ represents a spatial displacement, or shift with which the reflected particle leaves the scatterer, with $\xi(x^{\prime})$ giving the corresponding probability amplitude.
The usual dilemma is now evident. In the von Neumann example of Appendix D, interaction with the pointer alters the likelihood of the post-selection. Similarly, the reflection probability $P_{R}(t)\equiv\int dx|\Psi_{R}(x,t)|^{2}$ depends on the WP’s width $\Delta x$. An accurate measurement, $\Delta x\to 0$, destroys reflection, $P_{R}(t)\to 0$, since most of the WP broad in the momentum space, and is now transmitted. Conversely, the choice of a nearly monochromatic WP, $\Delta x\to\infty$, preserves $P_{R}(t)\to|R(p)|^{2}$ but fails to distinguish between different shifts $x^{\prime}$. A similar claim that, in a double slit experiment, one cannot both know the slit chosen by the particle and retain the interference pattern on the screen is often used to illustrate the Uncertainty Principle FeynL .
It is instructive to see how the notion of reflection time occurs in the classical limit of the just described “measurement”.
IV (Semi)classical reflection time
Consider the semiclassical limit of Eqs.(11) in the case of the barrier shown in Fig.1a. For energies less than the barrier height the reflection amplitude is given by Land
$R(k)\approx-i\exp\left[2i\int_{0}^{x_{<}(k)}q(x,k)dx\right],$ (14)
where $q(x,k)=\sqrt{k^{2}-2mV(x)}$ is the local momentum, and $x_{<}$ is the turning point, where $q(x_{<}(k),k)=0$ (see Fig.1a). The integral over $k$ in Eq.(3) contains rapidly oscillating exponentials and can be evaluated by the stationary phase method St . Integral over $dx^{\prime}$ in Eq.(8) is also oscillatory and, as is easy to check, has a stationary point at $x^{\prime}_{s}(p)$,
$\frac{mx_{s}^{\prime}(p)}{p}=2\int_{0}^{x_{<}(p)}\frac{mdx}{q(x,p)}.$ (15)
As a result a single envelope is selected in Eq.(11),
$\Psi_{R}(x,t)\approx\exp[-ip(x+x_{0})-iE(p)t]\quad\quad\quad\quad\quad\quad\quad$ (16)
$G_{0}(x_{s}^{\prime}(p)-X(t)).\quad$
We note that Eq.(15) equates the time it takes the free particle to travel a distance $x^{\prime}$ to the actual time of travel from the origin $x=0$ to the turning point $x_{<}$ and back. The classical time measurement now looks like this. The experimenter prepares a classical particle with momentum $p$ at $x_{0}<0$ to the left of the barrier, and waits until it bounces off it. Then he determines its position $x$ at time $t$, calculates $x^{\prime}_{s}=X(t)=x+x_{0}+vt$, and finally finds the duration spent in potential,
$\tau=\frac{mx_{s}^{\prime}}{p}\equiv\frac{x^{\prime}_{s}}{v}.$ (17)
For a semiclassical particle, prepared in a Gaussian state, a convenient reference point is the centre of mass of the reflected WP (16),
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{0}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{0}dx\,\Psi_{R}(x,t)|^{2}}$ (18)
from which one obtains
${\langle}\tau{\rangle}\equiv\frac{{\langle}x{\rangle}}{v}+\frac{x_{0}}{v}+t.$ (19)
The classical experiment cannot, however, determine such a duration for a reflected particle with energy higher than the barrier top, or for a particle reflected by a potential well. Such events are, however, possible in the quantum case which we will consider next.
V Quantum “reflection time”
One way to extend the classical result (17) to a fully quantum case is to associate a duration $\tau=x^{\prime}/v$ with each spatial shift $x^{\prime}$ in Eq.(11). Now having found the reflected particle in $x$ at $t$, one can assume that it had spent in the potential a duration $X(t)/v-\Delta x/v\lesssim\tau\lesssim X(t)/v+\Delta x/v$. The danger, as will be shown shortly, is that one will encounter negative durations whenever $V(x)$ provides negative shifts, $\xi(x^{\prime}<0)\neq 0$. We will proceed nonetheless.
For an accurate measurement, one requires $\Delta x\to 0$, and since $\xi(\tau)$ is a smooth function, one finds ($x<0$),
$|\Psi_{R}(x,t)|^{2}\approx\left|\int_{-\infty}^{\infty}dx^{\prime}G_{0}(x^{\prime})\right|^{2}\left|\xi(x+vt+x_{0})\right|^{2}.\quad$ (20)
With interference between the alternatives destroyed, a single duration $\tau(x)=t+\frac{x}{v}+\frac{x_{0}}{v}$ can be ascribed to a particle found in $x$. Averaging over all positions yields the mean “reflection time” (19)
${\langle}\tau{\rangle}{\xrightarrow[\Delta x\to 0]{}}\frac{1}{v}\frac{\int_{-\infty}^{\infty}dx\,x\,|\xi(x+vt+x_{0})|^{2}}{\int_{-\infty}^{\infty}dx\,|\xi(x+vt+x_{0})|^{2}}.$ (21)
The danger here, we repeat, is that, since $\xi(x^{\prime})$ may not vanish for $x^{\prime}<0$, the average (21) can, in principle, turn negative, and no longer represent a meaningful duration.
We note also that since $G_{0}(x){\xrightarrow[\Delta x\to 0]{}}(2\pi\Delta x^{2})^{1/4}\delta(x)$, the first integral in Eq.(20) behaves as $\sim\Delta x^{1/2}$, and $|\Psi_{R}(x,t)|^{2}\to 0$, because most of the momenta now go above the barrier (see also Sect.III).
Next we consider the case of a highly inaccurate measurement, $\Delta x\to\infty$, where all the alternatives, be they spatial shifts $x^{\prime}$, or times $x^{\prime}/v$, interfere to produce the probability $|\Psi_{R}(x,t)|^{2}$ of finding the reflected particle in $x$.
VI “Phase time” for reflection
As the initial wave packet becomes broader, $\Delta x\to\infty$, $G_{0}(x^{\prime}-X(t))$ in Eq.(8) varies little across the range of the $x^{\prime}$s contained in the $\xi_{\tau}(\tau)$, and can be approximated by the first two terms of its Taylor series, $G_{0}(x^{\prime}-X(t))\approx G_{0}(X(t))-x^{\prime}\partial_{x}G_{0}(X(t))$. Inserting this into (11), yields
$\Psi_{R}(x,t)\approx\exp[-ip(x+x_{0})-iE(p)t]\times\quad\quad\quad\quad\quad\quad\quad$ (22)
$G_{0}(X(t))\int dx^{\prime}\eta(p,x^{\prime})-\partial_{x}G_{0}(X(t))\int dx^{\prime}x^{\prime}\eta(p,x^{\prime})\quad$
where we wrote $\eta(p,x^{\prime})\equiv\exp(ipx^{\prime})\xi(x^{\prime})$. Inserting (22) into (18), and using (19) yields the “phase time”
${\langle}\tau{\rangle}{\xrightarrow[\Delta x\to\infty]{}}\frac{1}{v}\text{Re}\left[\frac{\int_{-\infty}^{\infty}dx^{\prime}x^{\prime}\eta(p,x^{\prime})}{\int_{-\infty}^{\infty}dx^{\prime}\eta(p,x^{\prime})}\right]$ (23)
$=\frac{1}{v}\frac{\partial\phi_{R}(p)}{\partial k}\equiv\tau_{ph},\quad\quad$
where $\phi_{R}(p)$ is the phase of the reflection coefficient, $R(p)=|R(p)|\exp[i\phi_{R}(p)]$. The phase time (23) is often thought to represent the duration a particle with momentum $p$ spends in the reflecting potential. [In quantum measurements, the complex quantity in the brackets is known as “weak value of $x^{\prime}$W1 , W2 . It can be shown that, as in the case of transmission W3 , its imaginary part determines the small change in the reflected particle’s mean momentum, which occurs because higher momenta are transmitted more easily, and are less likely to be reflected (see also GH4 ).]
However, one needs to be even more cautious. Since complex valued $\eta(p,x^{\prime})$ may change sign, $\tau_{ph}$ can be negative, even if all phase shifts in Eq.(13) are positive and $\xi(x^{\prime})$ vanishes for $x^{\prime}<0$ (for an example see Neg ).
Simple illustrations will be given next.
VII Zero-range potentials. Negative “reflection times”
For a particle scattered by a zero-range potential $V(x)=\Omega\delta(x)$, where $\delta(x)$ is the Dirac delta, the reflection amplitude is well-known to be
$R(k)=\frac{-im\Omega}{k+im\Omega}.$ (24)
Thus, we have
$\xi(x^{\prime})=m|\Omega|\exp(-m|\Omega x^{\prime}|)\,\theta(\pm x^{\prime}),$ (25)
where the upper and lower signs are for a barrier, $\Omega>0$, and a well $\Omega<0$, respectively.
For a $\delta$-barrier, both the average (21) and the phase time (23) are positive
${\langle}\tau{\rangle}{\xrightarrow[\Delta x\to 0]{}}\frac{1}{2|\Omega|v}>0,\quad\ \tau_{ph}(p)=v^{-1}\frac{m|\Omega|}{p^{2}+m^{2}\Omega^{2}}>0.\quad$ (26)
For a $\delta$-well FOOT both ${\langle}\tau{\rangle}$ and $\tau_{ph}$ change sign and become negative,
${\langle}\tau{\rangle}{\xrightarrow[\Delta x\to 0]{}}\frac{-1}{2|\Omega|v}<0,\quad\tau_{ph}(p)=v^{-1}\frac{-m|\Omega|}{p^{2}+m^{2}\Omega^{2}}<0.\quad$ (27)
The quantities in Eqs.(27) cannot be interpreted as durations the particle spends in the right half-space. The fault is with conversion of the phase shifts in Eq.(13) into temporal intervals, which yields a negative time $\tau$ whenever $x^{\prime}<0$. Finally, the reflection probability
$P_{R}{\xrightarrow[\Delta x\to 0]{}}\sqrt{2\pi}\Delta x\,m|\Omega|/2\quad\quad$ (28)
vanishes, as expected, for a WP narrow in the coordinate space.
VIII A step potential. Infinite phase time
Consider next a step potential
$V(x)=V\theta(x),$ (29)
for which the reflection amplitude is easily found to be
$R(k,q(k))=\frac{k-q}{k+q},$ (30)
where $q(k)=\sqrt{k^{2}-2mV}$. Evaluation of the integral (5) yields (for details see Appendix E)
$\xi(x^{\prime})=-\frac{2}{x^{\prime}}\,J_{2}(\sqrt{2mV}x^{\prime})\,\theta(x^{\prime})$ (31)
where $J_{2}(z)$ is the Bessel function of the first kind of order two Abram (see Fig. 2).
Figure 2: The amplitude distribution $\xi(x^{\prime})$ for a step potential (29), as given by (31).
It is readily seen that $\int_{0}^{\infty}dx^{\prime}\xi(x^{\prime})=-1$ and, as the height of the step increases, $\xi(x^{\prime}){\xrightarrow[V\to\infty]{}}-\delta(x^{\prime})$, and the reflection amplitude tends to that of a hard wall,
$R(p)=\int_{0}^{\infty}dx^{\prime}\exp(ipx^{\prime})\xi(x^{\prime}){\xrightarrow[V\to\infty]{}}-1.$ (32)
Figure 3: Integral $R(p,z)\equiv\int_{0}^{z}dx^{\prime}exp(ipx^{\prime})\xi(x^{\prime})$ converges to the reflection amplitude R(p)R(p)(dashed) as $z\to\infty$. Convergence is slower for energies close to the step’s height, $p\approx\sqrt{2mV}$
To understand the behaviour of the phase time $\tau_{ph}$, we note first that the integrand in the denominator of Eq.(23) contains the term $\exp[i(p-\sqrt{2mV})x^{\prime}]/x^{\prime 3/2}$, and converges, as illustrated in Fig. 3. The integrand in the numerator of Eq.(23), fall off as $\sim 1/x^{\prime 1/2}$, and
$\tau_{ph}(p)=\begin{cases}\frac{2m}{p\sqrt{2mV-p^{2}}}&\text{ for}\quad 0\leq p\leq\sqrt{2mV}\\ 0&\text{for}\quad p>\sqrt{2mV},\end{cases}$ (33)
becomes infinite as $p\to\sqrt{2mV}$. The fault is with the approximation (22), without which both ${\langle}x{\rangle}$ in Eq.(18) and the ${\langle}\tau{\rangle}$ defined by Eq.(21) remain finite, as shown in Fig.4 (see also GH4 ).
Figure 4: The centre-of-mass delay for a WP incident on the potential step (29) for two coordinate widths. Also shown by the dashed line is the phase time (23), which diverges when $E(p)=V$. The units are $\underline{x}=1/\sqrt{2mV}$ and $\underline{\tau}=p^{-1}(2V/m)^{-1/2}$.
IX Soft barrier as a superposition of hard walls
Having abandoned attempts to describe quantum reflection in terms of a “reflection time”, we look at the problem from a different angle. In Eq.(8) the free state $\Psi_{0}(x^{\prime}-x,t)$ differs only by sign from the state $\Psi_{R}^{\infty}(x,t|a)$, reflected by an infinite hard wall placed at $a=x^{\prime}/2$, and we can write
$\Psi_{R}(x,t)=\int_{0}^{\infty}da\,\xi_{a}(a)\Psi^{\infty}_{R}(x,t|a),\quad[x<0].$ (34)
where $\xi_{a}(a)=-2\;\xi(2a)$, and
$\Psi^{\infty}_{R}(x,t|a)=\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$ (35)
$-\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}\exp[ik(2a-x)-iE(k)t]\quad$
As long as only the reflected part of the WP is of interest, one can replace the actual potential $V(x)$ by a superposition of hard walls, $V_{\infty}(x|a)=0$ for $x<a$, and $\infty$ otherwise. The reflected WP is now given by the superposition of the states reflected by all infinite walls, with appropriate amplitude weights $\xi_{a}(a)$. Now $x^{\prime}=0$ corresponds to reflection by a wall placed at the origin, in which case the particle does not enter the region $x>0$. For a positive $x^{\prime}>0$ the wall is placed inside the $x>0$ region, so the centre of the WP, reflected from it, lags behind $\Psi^{\infty}_{R}(x,t|0)$. The opposite is true for $x^{\prime}<0$.
If the potential supports bound states, some of the walls will have to be placed in the $x<0$ region, suggesting reflections occurring even before the particle impacts the potential [cf. Eq.(27)]. However, for a pure barrier, having no poles in the upper half of the complex $k$-plane, we have
$\xi_{a}(a)\equiv 0,\quad\text{for}\quad a<0,$ (36)
so that all fictitious walls are found in the region $x>0$ (see Fig.1c). Next we use Eq.(34) to analyse reflection in two spatial dimensions.
Figure 5: a) In two dimensions, a Gaussian beam is incident on a potential barrier $V(x,y)=V(x)\theta(x)$. The reflected beam is a superposition of all beams, reflected by a variety of infinite walls placed at $x>0$. b) The centre of the reflected beam [cf. Eq.(44)] and the Goos-Hänchen shift (46).
X Quantum Goos-Hänchen effect
In two dimensions ${\bf r}=(x,y)$, ${{\bf k}}=(k_{x},k_{y})$], consider a potential occupying the $x>0$ half-space $V(x,y)=V(x)\theta(x)$. Now if the particle’s initial state, localised in the potential-free region $x<0$, is given by
$\Psi({\bf r},0)=\int_{-\infty}^{\infty}d{\bf k}{\langle}{\bf r}|{\bf k}{\rangle}{\langle}{\bf k}|\Psi(0){\rangle},\quad{\langle}{\bf r}|{\bf k}{\rangle}=\frac{\exp(i{\bf k}{\bf r})}{2\pi},\quad$ (37)
for the reflected WF one finds
$\Psi_{R}({\bf r},t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}dk_{x}\int_{-\infty}^{\infty}dk_{y}R(k_{x})\times$ (38)
$\exp[-ik_{x}x+ik_{y}y-iE(k)t]{\langle}{\bf k}|\Psi(0){\rangle}.$
where $R(k)$ is the reflection coefficient for one dimensional potential $V(x)$.
Acting as before, for a barrier not supporting bound states we find
$\Psi_{R}({\bf r},t)=\int_{0}^{\infty}da\,\xi_{a}(a)\,\Psi_{R}^{\infty}({\bf r},t|a)\quad[x<0],$ (39)
where $\Psi^{\infty}_{R}({\bf r},t|a)$ is the freely propagating initial state, reflected by a hard wall placed at $x=a$,
$\Psi^{\infty}_{R}(x,y,t|a)=\Psi_{0}(2a-x,y,t).$ (40)
Let $\Psi_{0}(x,y,t)$ be a transverse Gaussian beam (or a WP very broad in the direction of propagation), making an angle $\theta$ with the normal to the surface $x=0$ (see Fig. 5a). In the coordinate system $(x_{1},y_{2})$, obtained by rotating the axes $(x,y)$ counter clockwise by $\theta$, we have
$\Psi_{0}(x_{1},y_{1},t)=G_{0}(y_{1})\exp(ipx_{1}-ip^{2}t/2m),$ (41)
where we neglect the beam’s broadening by putting,
$G_{0}(y)=\left(\frac{2}{\pi\Delta y^{2}}\right)^{1/4}\exp\left(-\frac{y^{2}}{\Delta y^{2}}\right).$ (42)
(see Appendix G). Returning to the spatial shifts (13), $x^{\prime}=2a$, after a simple exercise in trigonometry, we find ($x<0$)
$\Psi_{R}(x_{2},y_{2},t){\xrightarrow[x_{2}\to\infty]{}}\exp[ipx_{2}-iE(p)t]\times$ (43)
$\int_{0}^{\infty}dx^{\prime}G_{0}(y_{2}-x^{\prime}\sin\theta)\exp(ip_{x}x^{\prime})\xi(x^{\prime})$
where the axes $(x_{2},y_{2})$ are obtained by reflecting the axes $(x_{1},y_{1})$ about the plane $x=0$ (see Fig. 2a).] Again we have a kind of quantum measurement. Detecting the particle at some $y_{2}$ one can deduce that it has experienced a spatial shift in the interval $(y_{2}-\Delta y_{2})/\sin\theta\lesssim x^{\prime}\lesssim(y_{2}+\Delta y_{2})/\sin\theta$. The measurement of $x^{\prime}$ is the more accurate the smaller is the ratio between the width of the beam $\Delta y_{2}$ and the sine of the angle of incidence $\theta$, $\Delta x^{\prime}=\Delta y_{2}/\sin\theta$.
The quantity of interest is the Goos-Hänchen shift $\delta y_{GH}$, i.e., the distance by which the reflected beam is displaced relative to the same beam reflected by a hard wall placed at $x=0$ (see Fig.2b). We note that superposition of the Gaussians in Eq.(41) does not have to be a Gaussian itself. However, one can always calculate the centre of the beam, the way it was done in Eq.(19)
${\langle}y_{2}{\rangle}=\frac{\int_{0}^{\infty}dy_{2}\,y_{2}\,|\Psi_{R}(x_{2}\to\infty,y_{2},t)|^{2}}{\int_{0}^{\infty}dy_{2}\,|\Psi_{R}(x_{2}\to\infty,y_{2},t)|^{2}},$ (44)
and from it obtain
$S_{GH}=\frac{{\langle}y_{2}{\rangle}}{\cos\theta}.$ (45)
In the limits of a narrow and broad beam, we find expressions similar to Eqs.(21) and (23), respectively,
$S_{GH}{\xrightarrow[\Delta y/\sin\theta\to 0]{}}\tan(\theta)\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}|\xi(x^{\prime})|^{2}}{\int_{0}^{\infty}dx^{\prime}|\xi(x^{\prime})|^{2}}$ (46)
and
$S_{GH}{\xrightarrow[\Delta y/\sin\theta\to\infty]{}}\tan(\theta)\text{Re}\left[\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}{\int_{0}^{\infty}dx^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}\right]\quad$ (47)
$=\tan(\theta)\frac{\partial\phi_{R}(k)}{\partial k}{\biggr{\rvert}}_{k=p_{x}=p\cos\theta}\quad\quad\quad\quad$
To calculate ${\langle}y_{2}{\rangle}$ for any $\Delta y$, we write $G_{0}(y)=\int\tilde{G}_{0}(q)\exp(iyq)dq$ and use Eqs. (6) and (87). The result is
${\langle}y_{2}{\rangle}\equiv{\text{Im}}\left[\frac{\int_{-\infty}^{\infty}dqF^{*}(q,p_{x},\theta)\partial_{q}F(q,p_{x},\theta)}{\int_{-\infty}^{\infty}dq|F(q,p_{x},\theta)|^{2}}\right],$ (48)
$F(q,p_{x},\theta)\equiv R(p_{x}-q\sin\theta)\tilde{G}_{0}(q).\quad\quad$
Finally, in the paraxial approximation of Appendix G, effects of beam’s broadening can be included by replacing $G_{0}(y)$ and $\tilde{G}_{0}(q)$ In Eqs.(41) and (48), by their coordinate-dependents versions (91) and (93). This approach is also useful beyond the paraxial approximation. For example, consider a doubly Gaussian wave packet, reflected by potential step (see section VIII) in the regime of total reflection (i.e when all energies in the wave packet are much lower than the step height). In this case, both the longitudinal delay and the lateral displacement can be directly related to the average depth at which the particle is reflected,
$\langle x_{2}\rangle=-x_{0}-\frac{p_{0}}{m}t+\langle x\rangle\cos\theta,\quad\quad\langle y_{2}\rangle=\langle x\rangle\sin\theta.$ (49)
The average depth at which the reflection occurs is related to the derivative of the phase [see (33)] of the reflection coefficient (30)
$\langle x\rangle\approx\iint dk_{x_{2}}dk_{y_{2}}{\frac{d\phi_{R}(k_{x})}{dk_{x}}\bigg{|}_{k_{x}=k_{x_{2}}\cos\theta-k_{y_{2}}\sin\theta}}$ (50)
$\times\frac{2}{\pi}\frac{\exp(-2k_{y_{2}}^{2}/\Delta k_{y_{2}}^{2})}{\Delta k_{y_{2}}}\frac{\exp[-2(k_{x_{2}}-k_{0})^{2}/\Delta k_{x_{2}}^{2}]}{\Delta k_{x_{2}}}.$
Here both the wave packet widths $\Delta k_{x_{2}}$ and $\Delta k_{y_{2}}$ must be sufficiently small, so that adding them to the average momentum $k_{0}$ keeps the corresponding energies significantly lower than the barrier height.
XI Conclusions
Above we attempted to give a most detailed description of the quantum Goos-Hänchen effect that an elementary theory can offer. For a classical particle with a known momentum launched at oblique angle at a repulsive potential, the distance $S_{GH}$ between the exit and entry point at the surface, where the potential vanishes, can be found by multiplying its lateral velocity $v_{II}$ by the duration $\tau$ its trajectory spends in the potential: $S_{GH}=v_{II}\tau$.
A straightforward extension of this analysis to the quantum case is, however, not possible. In one dimension, the quantum state of a particle with a known momentum $p=mv$ is a plane wave, and the moments when it enters and leaves the potential region simply cannot be defined. One notes, however, that the reflected state is a superposition (13) of multiple plane waves with phase shits corresponding to displacements, or shifts, $x^{\prime}$, induced by the potential. One notes also that in experiment where one prepares the particle in a wave packet (WP) state with a mean momentum $p$, and then observes it at $x$, the shifts are restricted to the range, determined by the WP’s coordinate width $\Delta x$. In the classical limit of this “quantum measurement”, the unique displacement $x^{\prime}_{s}$, with which the reflected particle leaves the potential, corresponds to the stationary region where rapid oscillations of the amplitude distribution $\xi(x^{\prime})$ slow down. This classical value allows one to deduce the duration $\tau$ that the particle’s trajectory spends in the potential as $x^{\prime}_{s}/v$.
There are two good reasons why one cannot continue to use the time argument beyond the classical limit, e.g., by converting each spatial shift into a duration using $\tau(x^{\prime})=x^{\prime}/v$. One is that some of these “durations” will have to be negative in the case of a potential well, or, indeed, of any potential supporting bound states. The idea that a particle which which makes a U-turn before entering the potential spends in the scatterer a “negative duration” does not help explain the phenomenon in classical-like terms. [We are aware of the recently reported “experimental evidence of negative times” St , and remain sceptical.]. The notion that a particle’s state can acquire a negative phase shift can, however, be made without causing a further controversy.
The other reason is that, as with any quantum measurement of this kind, one faces a dilemma rooted in the Uncertainty Principle FeynL . One can determine the shift experienced by the particle accurately, but the reflection probability will not correspond to that of a particle with a momentum $p$. In fact, it will vanish. Or, one can keep the probability (almost) intact, but loose all the knowledge about the $x^{\prime}$s to interference. In this case it is still possible to find the displacement of the centre of mass of a very broad reflected WP and, from it the “phase time” (23). However, both are related to the first moment of an oscillatory complex valued distribution, and can be negative even for a potential barrier, where all shifts $x^{\prime}$s are positive (for an example see Neg ).
One can represent the scattering potential by a superposition of virtual hard walls, distributed in such a manner that the outgoing WP is a sum of all WPs reflected by each of wall. Such a representation is not without its problems, since in the case of a well, some of the fictitious walls would need to be placed in the potential-free region, implying that the scattering occurs even before the particle reaches the scatterer. Still the idea is somewhat less confusing than the one invoking negative durations, and can be used for illustrative purposes.
For example, the hard wall picture helps one to visualise what happens in the quantum Goos-Hänchen case GH1 . The outgoing Gaussian beam is composed of the beams reflected by each fictitious wall (see Fig. 5.a). The cross section of the beam’s wave function is a superposition of Gaussians (we neglect the lateral spreading of the wavepacket during propagation) similar to one would obtain in a von Neumann quantum measurement. As the width of the incident beam increases (the measurement becomes less accurate) the Goos-Hänchen shift changes from Eq.(46) to Eq.(61), involving the “strong” and the “weak” values of the coordinate shift $x^{\prime}$.
Appendix A. Scattering by a one-dimensional barrier
Consider scattering by a one-dimensional potential barrier $V(x)$, contained between $x=0$ and $x=d$,
$V(x)\equiv 0,\quad\mathrm{for}\quad x<0\quad\mathrm{and}\quad x>d.$ (51)
The left (right) scattering states, for a particle incident from the left (right),
${\langle}{x}|\Phi_{l}(k){\rangle}=\frac{1}{\sqrt{2\pi}}\left[e^{ikx}+R_{l}(k)e^{-ikx}\right],\quad x<0\quad$ (52)
$=\frac{1}{\sqrt{2\pi}}T_{l}(k)e^{ikx},\quad x>d\quad$
[…]
$=\frac{1}{\sqrt{2\pi}}T_{r}(k)e^{-ikx},x<0$
form a complete orthonormal basis,
$\int_{0}^{\infty}dk\left[|\Phi_{l}(k){\rangle}{\langle}\Phi_{l}(k)|+|\Phi_{r}(k){\rangle}{\langle}\Phi_{r}(k)|\right]=1.$ (53)
It follows from the Wronskian relations between the states (52) and the fact that the complex conjugate $\Phi^{*}_{l,r}(k,x)$ are also solutions of the Schrödinger equation (SE), that
$T_{l}(k)=T_{r}(k)=T(k),\quad R_{l}(k)T^{*}(k)=-R^{*}_{r}(k)T(k),\quad\quad$ (54)
$T(-k^{*})=T(k)^{*},\quad R_{l.r}(-k^{*})=R^{*}_{l.r}(k).\quad\quad\quad\quad\quad\quad\quad$
The initial state is a wave packet $|\Psi(x,0){\rangle}$ completely localised to the left of the barrier, which we expand in plane waves $|k{\rangle}$ as
$|\Psi(x,0){\rangle}=\int_{-\infty}^{\infty}dk|k{\rangle}{\langle}k|\Psi(0){\rangle}\equiv\int_{-\infty}^{\infty}dkA(k)|k{\rangle},$ (55)
${\langle}x|k{\rangle}=\frac{1}{\sqrt{2\pi}}e^{ikx},$
To evolve the state in time we expand it in the scattering states (52),
$\Psi(x,t)=\int_{0}^{\infty}dk{\langle}\Phi_{r}(k)|\Psi(0){\rangle}e^{-iE(k)t}{\langle}x|\Phi_{r}(k){\rangle}+$ (56)
$\int_{0}^{\infty}dk{\langle}\Phi_{l}(k)|\Psi(0){\rangle}e^{-iE(k)t}{\langle}x|\Phi_{l}(k){\rangle},$
where $E(k)=k^{2}/2m$, and $m$ is the particle’s mass. Using $\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}|x^{\prime}{\rangle}{\langle}x^{\prime}|k^{\prime}{\rangle}{\langle}k^{\prime}|=1$ we find
${\langle}\Phi_{l,r}(k)|\Psi(0){\rangle}=\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}$ (57)
${\langle}\Phi_{l,r}(k)|x^{\prime}{\rangle}{\langle}x^{\prime}|k^{\prime}{\rangle}{\langle}k^{\prime}|\Psi(0){\rangle}$
Now, since the initial state lies to the left of the barrier, i.e., at $x<0$, we can use the corresponding expressions for $\Phi^{*}_{l,r}(k,x)$ [cf. Eq.(52)] and extend integration over $x^{\prime}$ to $(-\infty,\infty)$. With $(2\pi)^{-1}\int_{-\infty}^{\infty}\exp(ikx)\,dx=\delta(k)$, this gives
${\langle}\Phi_{l}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}\quad\quad\quad$ (58)
$\left[e^{-ikx^{\prime}}+R_{l}^{*}(k)e^{ikx^{\prime}}\right]e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=\quad\quad\quad\quad\quad$
${\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle},\quad\quad\quad\quad\quad\quad\quad\quad$
${\langle}\Phi_{r}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}\quad\quad\quad$
$T^{*}(k)e^{ikx^{\prime}}e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=T^{*}(k){\langle}-k|\Psi(0){\rangle}.\$
[…]
$\Psi(x,t)=\frac{1}{\sqrt{2\pi}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$ (59)
$\{\int_{0}^{\infty}dk[{\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle}]T(k)e^{ikx-iE(k)t}$
$+\int_{0}^{\infty}dkT^{*}(k){\langle}-k|\Psi(0){\rangle}[e^{-ikx}+R_{r}(k)e^{ikx}]e^{-iE(k)t}\},$
which, after taking into account (54), and changing $k\to-k$ simplifies to
$\Psi(x,t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dkT(k){\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}.$ (60)
[…]
$\Psi(x,p,t)=\int_{0}^{\infty}dk[{\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle}]$ (61)
$\times\frac{1}{\sqrt{2\pi}}[e^{ikx}+R_{l}(k)e^{-ikx}]e^{-iE(k)t}\quad\quad\quad\quad$
$+\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}dkT^{*}(k){\langle}-k|\Psi(0){\rangle}T(k)e^{-ikx}e^{-iE(k)t},$
Or, recalling that $|T(k)|^{2}+|R_{l}(k)|^{2}=1$ and $R_{l}^{*}(k)=R_{l}(-k)$,
$\Psi(x,t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}+$ (62)
$\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}R_{l}(k)dk{\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}.$
Validity of (62) is easily proven. Define
$I_{1}=\quad\int_{0}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}$ (63)
$I_{2}=\int_{0}^{\infty}dkR_{l}^{*}(k){\langle}-k|\Psi(0){\rangle}e^{ikx-iE(k)t}+$ (64)
$\int_{0}^{\infty}dkR_{l}(k){\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}=$
$\int_{-\infty}^{\infty}dkR_{l}(k){\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}$
$I_{3}=\int_{0}^{\infty}dk|R_{l}(k)|^{2}{\langle}-k|\Psi(0){\rangle}e^{-ikx-iE(k)t}+$ (65)
$\int_{0}^{\infty}dk|T(k)|^{2}{\langle}-k|\Psi(0){\rangle}e^{-ikx-iE(k)t}=$
$\int_{0}^{\infty}dk{\langle}-k|\Psi(0){\rangle}e^{-ikx-iE(k)t}=$
$\int_{-\infty}^{0}dk{\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}$
Then adding $I_{1}+I_{2}+I_{3}$ yields $\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}$
Equations (60) and (62) allow one to monitor the time evolution of the wave function in the regions $x<0$ and $x>d$, but not inside the potential, $0<x<d$, where one needs to know the behaviour of $\Phi_{l.r}(k,x)$, not specified by Eqs.(52). For example, the initial WP located to the left of the potential, is represented by the first term in Eq.(62). As time progresses, it gradually “disappears from the screen”, and is replaced by the second term in (62) in the region $x<0$, while the transmitted part (60) appears in the $x>0$ region. Note that (60) and (61) are valid even when the WP expansion (67) contains negative momenta, propagating away from the barrier, as we illustrate next.
Appendix B. Gaussian wave packets
It is common practice to study scattering of Gaussian states. Choosing
${\langle}k|\Psi(0){\rangle}=\left({2\pi^{-1}\Delta k^{2}}\right)^{-1/4}\exp\left[-\frac{(k-p)^{2}}{\Delta k^{2}}-ikx_{0}\right]\quad$ (66)
with $x_{0}<0$ and $|x_{0}|>>2/\Delta k$ yields an initial Gaussian wave packet, set a distant $|x_{0}|$ to the left of the barrier, and moving with velocity $v=p/m$ towards it or away from it, depending on the sign of $p$
${\langle}x|\Psi(0){\rangle}=\left(\frac{\pi\Delta x^{2}}{2}\right)^{-1/4}\exp\left[-\frac{(x-x_{0})^{2}}{\Delta x^{2}}+ip(x-x_{0})\right]\quad$ (67)
where $\Delta x=2/\Delta k$. For $p<\Delta k$, the initial WP contains also negative momenta, even though its move towards the barrier. Similarly, with $p<0$, the WP moving now away from the barrier contains positive momenta. The first term in Eq.(62), describing free motion of the initial wave packet, is now given by
$\Psi(x,t)=\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{ikx-iE(k)t}$ (68)
$=e^{ip(x-x_{0})-iE(p)t}G_{0}(x-pt/m,t)$
where the envelope $G$, which contains the effects of spreading is given by
$G_{0}(x-\frac{pt}{m},t)=\left(\frac{2\Delta x^{2}}{\pi\Delta x_{t}^{4}}\right)^{1/4}\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x_{t}^{2}}\right],\quad$ (69)
$\Delta x_{t}\equiv\sqrt{\Delta x^{2}+\frac{2it}{m}},\quad\quad\quad\quad\quad\quad\quad$
If the spreading can be neglected, e.g., for $2t/m<<\Delta x^{2}$, the envelope travels without distortion with a velocity $v=p/m$,
$G_{0}(x-\frac{pt}{m},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(x-vt-x_{0})^{2}}{\Delta x^{2}}\right].\quad$ (70)
Equations (68)-(70) are valid, we repeat if the initial Gaussian state (66) contains momenta in the direction opposite to the WP’s velocity $v$.
Appendix C. Free motion and reflection from an infinite (hard) wall
To check the consistency of Eqs.(60) and (62), we consider two simple cases. For free motion, $V(x)\equiv 0$, $T(k)=1$, $R(k)=0$ for $\infty<k<\infty$, Eq.(60) recovers, as expected, the correct answer (68) for $x>d$.
More interesting is the case of reflection by an infinite wall placed at some $0<a$, $V(x)=0$ for $x<a$, and $V(x)=\infty$ for $x\geq a$. Now we have
$R(k)=-\exp(2ika),\quad-\infty<k<\infty,$ (71)
$\xi(x^{\prime})=-\delta(x^{\prime}-2a),\quad\quad\quad\quad\quad\quad\quad$
where the negative sign ensures that the wave function vanishes at $x=a$. Now the reflected WP entering the region $x<0$ is, apart from the change of the sign, a copy of the freely propagating initial WP, reflected about $x=a$.
$\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dkR_{l}(k){\langle}k|\Psi(0){\rangle}e^{-ikx-iE(k)t}=$ (72)
$-\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dk{\langle}k|\Psi(0){\rangle}e^{-ik(x-2a)-iE(k)t}.$
This is also the correct result which we will need in the following.
Appendix D. Quantum measurements
Consider, in one dimension, a particle (pointer) (P) with position $f$ and momentum $\lambda$}, briefly interacting with a quantum system (S),
$\hat{H}=\hat{H}^{S}+\hat{H}^{P}+\hat{H}_{int},$ (73)
where
$\hat{H}^{P}=\hat{\lambda}^{2}/2\mu,\quad\hat{H}_{int}=\hat{\lambda}\hat{B}\delta(t+0),$ (74)
where $\mu$ is the pointer’s mass, and $\hat{B}=\sum_{i}|b_{i}{\rangle}B_{i}{\langle}b_{i}|$ is the system’s operator to be measured. The system is pre- and post-selected in states $|I{\rangle}$ and $|F{\rangle}$ at $t=0$ and $t$, respectively. At $t=0$ the pointer is prepared in the Gaussian state (67) $x\to f$, $p\to{\lambda}$, $\Delta x\to\Delta f$, $x_{0}\to f_{0}=0$]. At time $t$ the observer accurately determines the pointer’s position, hoping in this way to learn something about the value of $\hat{B}$ at the time the system interacted with the pointer. The pointer’s state after the system was found in $|F{\rangle}$ at $t$, is
${\langle}F|\Phi(t){\rangle}=\sum_{i}\Psi(f-B_{i},t)\mathcal{A}^{S}(F\leftarrow b_{i}\leftarrow I),$ (75)
where $\Psi(f-B_{i},t)$ is given by Eqs.(68)-(69), and $\mathcal{A}_{i}^{S}$ is the system’s transition amplitude,
$\mathcal{A}^{S}(F\leftarrow b_{i}\leftarrow I)\equiv{\langle}F|\exp(-i\hat{H}^{S}t)|b_{i}{\rangle}{\langle}b_{i}|I{\rangle}.$ (76)
It is convenient to introduce an amplitude distribution $\xi(f^{\prime})$ (if required, the sum can be replaced by an integral)
$\xi(f^{\prime})\equiv\sum_{i}\delta(f^{\prime}-B_{i})\mathcal{A}^{S}(F\leftarrow b_{i}\leftarrow I),$ (77)
so that (75) can be rewritten as
$\Phi(f,t)\equiv{\langle}f|\Phi(t){\rangle}=\int df^{\prime}\Psi(f-f^{\prime},t)\xi(f^{\prime}),$ (78)
where $\xi(f^{\prime})$ is the probability amplitude of the quantity, represented by the operator $\hat{B}$, having a value $f^{\prime}$. Suppose we can neglect the spreading of the pointer’s wave packet, prepared with mean momentum ${\lambda}=0$, $\Psi(f-f^{\prime},t)\approx\left(\frac{2}{\pi\Delta x^{2}}\right)^{1/4}\exp\left[-\frac{(f-f^{\prime})^{2}}{\Delta f^{2}}\right]\equiv G_{0}(f-f^{\prime})$. Now only the values in the interval $f-\Delta f\lesssim f^{\prime}\lesssim f+\Delta f$ contribute to $\Phi(f,t)$ and by determining $f$ accurately one measures $\hat{B}$ to a (quantum) accuracy $\Delta f$. Note that as $\Delta f\to\infty$ the probability of the post-selection, $P(F)$, is no longer affected by the pointer, $P(F)\to|{\langle}F|\exp(-i\hat{H}^{S}t|I{\rangle}|^{2}$.
If the spreading can still be neglected, but the pointer is prepared with non-zero mean momentum ${\lambda}\neq 0$,
$\Phi(f,t)\equiv{\langle}f|\Phi(t){\rangle}=\exp[i\lambda f-iE(\lambda)t]\times$ (79)
$\int df^{\prime}G_{0}(f-f^{\prime})\exp(-i\lambda f^{\prime})\xi(f^{\prime}),$
and finding the pointer at $f$ one can still know that $\hat{B}$ had the value roughly in the interval $(f-\Delta f,f+\Delta f)$. However, now sending $\Delta f\to\infty$ does not fully decouple the pointer from the system, as $P(F)\to|\int df^{\prime}\exp(-i{\lambda}f^{\prime})\xi(f^{\prime})|^{2}\neq|{\langle}F|\exp(-i\hat{H}^{S}t|I{\rangle}$. The similarity between Eqs.(78) and (8) provides the required analogy.
Appendix E. Rectangular barrier and potential step
Consider a rectangular barrier $V(x)$ of a hight $V=const$, occupying the region $0\leq x\leq d$. For a particle of mass $m$, the reflection amplitude is well known to be
$R(k)=\frac{(k^{2}-q^{2})\sin qd}{2ikq\cos qd+(k^{2}+q^{2})\sin qd}$ (80)
where $q(k)\equiv\sqrt{k^{2}-2mV}\equiv\sqrt{k^{2}-W^{2}}$. The amplitude is an entire function in the complex $k$-plane, unchanged by replacement $q\to-q$ and satisfying, for real $k$ , $R(-k)=R^{*}(k)$. It has poles in the lower half plane [complex zeroes of the denominator in Eq.(80)], associated with the resonances at the energy above the barrier, $k^{2}/2m>V$. Thus to calculate $\xi(x^{\prime})=(2\pi)^{-1}\int_{-\infty}^{\infty}dk\exp(-ikx^{\prime})R(k)$ we may shift the contour of integration into the upper half-plane by putting $k=k+i\epsilon$. $\epsilon\to 0$, so $\xi(x^{\prime})=0$ for $x^{\prime}<0$. Note that on the contour shown in Fig.6 q(k)=q(k)q(-k)=q(k) for $|k|\leq W$, and q(k)=q(k)q(-k)=-q(k) for $|k|>W$,
To obtain the reflection amplitude for a potential step, $V(x)=V\theta(x)$, we send $d\to\infty$ before taking the limit $\epsilon\to 0$, and retain only the growing exponentials. (For example, for $k>W$, we have $q(k+i\epsilon)\approx q(k)+ik\epsilon/q$ so that $\exp[-iq(k+i\epsilon)d]\to\infty$ as $d\to\infty$.) This gives the well known expression
$R(k,q(k))=\frac{k-q}{k+q},$ (81)
which has no poles, but two branch cuts from -$\infty$ to W-W, and from $W$ to $\infty$, formed by the coalescence of the resonance poles as $d$ increases. To evaluate the amplitude distribution of the shifts, $\xi(x^{\prime})$, we integrate along the contour passing just above the cuts (see Fig.6). For $x^{\prime}<0$ the contour can be closed in the upper half-plane where $R(k)$ has no singularities, so $\xi(x^{\prime})\equiv 0$.
Figure 6: For $x^{\prime}>0$, the contour of integration running just above the real $k$-axis can be replaced by a loop around the branching point at $k=\pm\sqrt{2mV}$
For $x^{\prime}>0$, the contour can be transformed into a loop around the branch points as shown in Fig.4, and we have
$\xi(x^{\prime}\geq 0)=(2\pi)^{-1}\oint dkR(k)\exp(-ikx^{\prime})\quad\quad$ (82)
$=-\frac{4W}{\pi}\int_{0}^{1}dtt\sqrt{1-t^{2}}\sin(Wx^{\prime}t),\quad\quad\quad\quad$
$=\frac{2}{\sqrt{2mV}}\frac{d}{dx^{\prime}}\left(\frac{J_{1}(\sqrt{2mV}x^{\prime})}{x^{\prime}}\right),\quad\quad\quad\quad$
and $\xi(x^{\prime}<0)\equiv 0$. In (82) $J_{1}(z)$ is the Bessel function of the first order Abram . It is easy to check that $\int_{0}^{\infty}\xi(x^{\prime})=R(p=0)=-1$. Fimally, using Eqs. 9.1.30 of Abram , one obtains
$\xi(x^{\prime}\geq 0)=-\frac{2}{x^{\prime}}\,J_{2}(\sqrt{2mV}x^{\prime}).$ (83)
Appendix F. Centre-of-mass of the reflected wave packet
Calculation of the average
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{\infty}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{\infty}dx\,|\Psi_{R}(x,t)|^{2}}$ (84)
where
$\Psi_{R}(x,t)=\int_{-\infty}^{\infty}dkF(k,t)\exp[-ikx],$ (85)
$F(k,t)\equiv R_{l}(k){\langle}k|\Psi(0){\rangle}\exp[-iE(k)t].$
reduces to evaluation of simple quadratures. Insertion of the relations
$\int dx\exp[i(k-k^{\prime})x]=2\pi\delta(k-k^{\prime}),\quad\quad\quad\quad$ (86)
$\int dx\,x\,\exp[i(k-k^{\prime})x]=i\pi(\partial_{k^{\prime}}-\partial_{k})\delta(k-k^{\prime}).$
into (84)-(85) yields Eq.(48),
${\langle}x{\rangle}\equiv\frac{{\text{Im}}\left[\int_{-\infty}^{\infty}dkF^{*}(k,t)\partial_{k}F(k,t)\right]}{\int_{-\infty}^{\infty}dk|F(k,t)|^{2}}.$ (87)
Appendix G. Gaussian beam in paraxial approximation
Consider a monoenergetic two dimensional Gaussian beam, propagating freely along the $x$-axis,
$\Psi_{0}(x,y,t)=\exp(-ip^{2}t/2m)\times,\quad\quad\quad\quad$ (88)
$\int dq\tilde{G}_{0}(q)\exp[iqy+i\sqrt{p^{2}-q^{2}}(x-x_{0})],\quad$
where
$\tilde{G}_{0}(q)=({2\pi^{-1}\Delta q^{2}})^{-1/4}\exp(-q^{2}/\Delta q^{2}).$ (89)
If the spread of the tranverse moment $\Delta q$ is small compared to $p$, for $x>x_{0}$ we have
$\Psi_{0}(x,y,t)=\exp(ip(x-x_{0})-ip^{2}t/2m)\times$ (90)
$\int dq\tilde{G}_{0}(q,T)\exp(iqy)$
here $T=T(x)=(x-x_{0})/v$ is the time it takes particle with velocity $v=p/m$ to travel a distance $x-x_{0}$, and
$\tilde{G}_{0}(q,T)=({2\pi^{3}\Delta q^{2}})^{-1/4}\exp\left(-\frac{q^{2}}{\Delta q^{2}}-i\frac{q^{2}T}{2m}\right)$ (91)
Integration over $dq$ the yields the paraxial approximation for the beam
$\Psi_{0}(x,y,t)=\exp(ipx-ip^{2}t/2m)G_{0}(y,T(x))$ (92)
where the envelope
$G_{0}(y,T)=\left(\frac{2\Delta y^{2}}{\pi\Delta y_{T}^{4}}\right)^{1/4}\exp\left[\frac{-y^{2}}{\Delta y_{T}^{2}}\right],\quad\quad$ (93)
$\Delta y_{T}\equiv\sqrt{\Delta y^{2}+\frac{2iT}{m}}.\quad\quad\quad\quad\quad\quad\quad$
now accounts for beam’s broadening. Neglecting the broadening yields yields Eq.(41).
Acknowledgements
DS acknowledges financial support by the Grant PID2021-126273NB-I00 funded by MICINN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, as well as by the Basque Government Grant No. IT1470-22. YC is a research associate of the Fund for Scientific Research (F.R.S.-FNRS). YC acknowledges the support of the EU from the EIC Pathfinder Challenges 2022 call through the Research Grant 101115149 (project ARTEMIS). The content reflects only the authors’ views, and the European Commission is not responsible for any use that may be made of the information it contains.